find slope (equation)
m= -6; (-8,6)

Answer by uday1435(57)

The general equation is y = mx +b
pluging in x = -8, y = 6 and m = -6 we get
6 = (- 6)(-8) + b
6 = 48 + b
b = -42
now plugging in values of m ad b in the general equatio we get
y = - 6 x - 42
That is
6x + y + 42 = 0 is the required equation